ABC Company uses a Materials Inventory account to record both direct and indirect materials. ABC charges direct materials to$80,000. $90,000. $30,000. $50,000. $110,000.Maple Mount Fishery is a canning company in Astoria. The company uses a normal costing system in which factory overhead is ap$372,840. $375,000. $383,000. $386,000. $383,300.
Answer #1

Part 1

Total materials purchased = $90,000

Total materials issued to production = $110,000

Beginning materials inventory = $50,000

Total materials issued to production = Beginning materials inventory + Total materials issued to production- Ending materials inventory

110,000 = 50,000+90,000- Ending materials inventory

Ending materials inventory = 140,000-110,000

= $30,000

The ending materials inventory cost is $30,000.

Third option is correct.

Part 2

Raw materials, beginning = $61,600

Raw materials, ending = 61,600 - 61,600 x 10%

= 61,600-6,160

= $55,440

Direct materials used = $379,000

Direct materials purchased = ?

Direct materials used = Raw materials, beginning + Direct materials purchased - Raw materials, ending

379,000 = 61,600+Direct materials purchased - 55,440

Direct materials purchased = 379,000+55,440-61,600

= $372,840

The amount of direct materials purchased during the year is $372,840.

First option is correct.

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks